given TC=70+30Q+2Q2. P=40 find the optimalquantity produced and sold by the firms so that they are maximizing profits


1
Expert's answer
2021-12-13T16:52:36-0500

TC=70+30Q+2Q2. P=40

MC="\frac{\delta TC}{\delta Q}"= 30+4Q

MC=P

30+4Q=40

4Q=10

Q=2.5